# STRAP: Robot Sub-Trajectory Retrieval for Augmented Policy Learning

### v1

@@ 3.1 Dynamic Time Warping
A cumulative distance matrix $D$ is computed via dynamic programming as follows: $D(0,0)=C(0,0),D(n,1)=\sum_{k=1}^{n}C(k,1)\text{ for }n\in[1:N]\text{ and }D(1,m)=\sum_{k=1}^{m}C(1,k)\text{ for }m\in[1:M]$. Then the following dynamic programming calculation is performed:
@@
where $C(x_{i},y_{j})$ is the distance between points $x_{i}$ and $y_{j}$. We assume this cost matrix is pre-provided, and we describe how we compute this from raw camera images in Sec. 4.3. The optimal alignment between the sequences is found by backtracking from $D(n,m)$ to $D(0,0)$. This guarantees that the start is matched to the start and the end is matched to the end or that the pairs $(x_{0},y_{0})$ and $(x_{n},y_{m})$ are the start and end of the path. This optimal paring path consists of the best possible alignment between $X$ and $Y$ such that the cumulative cost between all matched pairs is minimized. DTW, as described, is widely used in time-series analysis, speech recognition, and other domains where temporal variations exist between sequences. In the context of our retrieval problem, DTW is used to go beyond retrieving exactly matched sequences to matching variable length subsequences, as we describe below.
@@ 4.1 Problem Setting: Retrieval-augmented Policy Learning
Given $\mathcal{D}_{\text{prior}}$ and $\mathcal{D}_{\text{target}}$, the goal is to learn a language-conditioned policy $\pi_{\theta}(a|s,l)$ that can predict optimal actions $a$ in the target environment when prompted with the current state $s$ and language instruction $l$. Assuming we can obtain a measure of success (such as task completion), and a broad set of initial conditions $s_{0}\sim\rho_{\text{test}}(s_{0})$ in the test environment. The objective of policy learning is to determine the policy parameters $\theta$ to maximize the expected success metric when evaluated on test conditions, under the policy $\pi_{\theta}$ and test-time environment dynamics. Since we are only provided a limited corpus of data, $\mathcal{D}_{\text{target}}$, in the target domain, these policy parameters cannot be learned by simply performing maximum likelihood on $\mathcal{D}_{\text{target}}$. Instead, we will present an approach where a smaller, “relevant” subset of the offline dataset $\mathcal{D}_{\text{retrieval}}$subseteq $\mathcal{D}_{\text{prior}}$ is retrieved non-parametrically and then mixed with the smaller in-domain dataset $\mathcal{D}_{\text{target}}$ to construct a larger, augmented training dataset, i.e., $\mathcal{D}_{\text{target}}\cup\mathcal{D}_{\text{retrieval}}$, which is most relevant to the desired test-time conditions $\rho_{\text{test}}(s_{0})$. This can then be used for training policies via imitation learning, as we will describe in Sec. 4.5. Doing so avoids an expensive generalist training procedure and rather focuses the learned model to being a high-performing specialist in a particular setting of interest. The key questions becomes - How can we define what subset of the offline dataset $\mathcal{D}_{\text{prior}}$ is relevant to construct $\mathcal{D}_{\text{retrieval}}$?
@@ 4.3 Foundation Model-Driven Relevance Metrics for Retrieval
In this work, we will adopt the insight that vision(-language) foundation models Oquab et al. (2023); Radford et al. (2021) offer off-the-shelf solutions to this problem of measuring the semantic and visual similarities between sub-trajectories, capturing object- and task-centric affordances, while being robust to low-level variations in scene appearance. Trained on web-scale real-world image(-text) data, these models are typically robust to low-level perceptual variations, while providing semantically rich representations that naturally capture a notion of object-ness and semantic correspondence. Denoting a vision foundation model as $\mathcal{F}(\cdot)$, we can compute the pairwise distance of two camera views with an L2 norm [^2] Other cost metrics such as (1-cosine similarity) could be used here as well. in embedding space, i.e., $||\mathcal{F}(o_{i})-\mathcal{F}(o_{j})||_{2}$. While aggregation methods such as temporal averaging could be used to go from embedding of a single image to that of a sub-trajectory, they lose out on the actions and dynamics. We instead opt for a sub-trajectory matching procedure based on the idea of DTW Giorgino (2009) and use the embeddings for finding maximally relevant sub-trajectories. Given two sub-trajectories, $\bm{t}_{i}$ and $\bm{t}_{j}$, we compute a pairwise cost matrix $C\in\mathbb{R}^{|\bm{t}_{i}|\times|\bm{t}_{j}|}$, where its value is as computed by:
$C(i,j)=||\mathcal{F}(o_{i})-\mathcal{F}(o_{j})||_{2}$ (2)
@@
3:$\mathcal{T}_{\text{target}}\leftarrow\texttt{SubTrajSegmentation}(\mathcal{D}_{\text{target}},\epsilon);$triangleright Heuristic demo chunking
4:$\mathcal{E}_{\text{prior}}\leftarrow\{\{\mathcal{F}(o_{t})|o_{t}\in T\}|T\in\mathcal{D}_{\text{prior}}\};$triangleright Embed $\mathcal{D}_{\text{prior}}$
5:$\mathcal{E}_{\text{target}}\leftarrow\{\{\mathcal{F}(o_{t})|o_{t}\in T\}|T\in\mathcal{T}_{\text{target}}\};$triangleright Embed chunked $\mathcal{D}_{\text{target}}$
@@
8:     $\mathcal{M}\leftarrow[];$triangleright Initialize empty match storage
@@
10:         $D\leftarrow\texttt{computeCostMatrix}(\mathcal{E}_{\text{target}},\mathcal{E}_{\text{prior}});$triangleright Eq. 2
11:         $\mathcal{M}_{i,j}\leftarrow\texttt{extractSubTrajectory}(D,T_{\text{prior}});$triangleright Dynamic Programming
@@
14:$\mathcal{D}_{\text{retrieval}}\leftarrow\texttt{retrieveTopKMatches}(\mathcal{M},K);$triangleright Sec. 4.4
@@
17:     sample $\mathcal{B}\sim\mathcal{D}_{\text{target}}\cup\mathcal{D}_{\text{retrieval}}$ to update policy $\pi_{\theta}$ with loss $\mathcal{L}(\mathcal{B};\theta)$triangleright Eq. 3
@@ 4.5 Putting it all together: STRAP
To start the retrieval process, we encode image observations in $\mathcal{D}_{\text{target}}$ and $\mathcal{D}_{\text{prior}}$ using a vision foundation model, e.g., DINOv2 Oquab et al. (2023) or CLIP Radford et al. (2021). To best leverage the multi-task trajectories in $\mathcal{D}_{\text{prior}}$, we split the demonstrations in $\mathcal{D}_{\text{target}}$ into atomic chunks based on a low-level motion heuristic. Then we generate matches between chunked $\mathcal{D}_{\text{target}}$ and $\mathcal{D}_{\text{prior}}$ and construct $\mathcal{D}_{\text{retrieval}}$ by selecting the top $K$ matches uniformly across all chunks. Combining $\mathcal{D}_{\text{retrieval}}$ with $\mathcal{D}_{\text{target}}$ forms our dataset for learning a policy. In a standard policy learning setting, noisy retrieval data can lead to negative transfer, e.g., when observations similar to the target data are labeled with actions that achieve a different task. Without conditioning, such contaminated samples hurt the policy’s downstream performance. We propose to use a language-conditioned policy to deal with this inconsistency. With conditioning, the policy can distinguish between samples from different tasks, separating misleading from expert actions while benefiting from positive transfer from the additional training data and context of the language conditioning.
@@ 5.1 Experimental Setup
  • DROID-Kitchen: Scaling STRAP to more realistic scenarios, we evaluate STRAP on vegetable pick-and-place in three real kitchen environments. We collect 50 multi-task demonstrations in every scene totaling 150 demonstrations in $\mathcal{D}_{\text{prior}}$ (Kitchen). Task combinations are sampled randomly from three possible tasks per environment and without replacement. Every demonstration consists of two unique tasks with randomized object poses and appearances. $\mathcal{D}_{\text{target}}$ is specific to each environment and contains 3 demonstrations of one of the three possible tasks. To investigate STRAP’s scalability to larger datasets, we construct an additional $\mathcal{D}_{\text{prior}}$ consisting of 5000 demonstrations from the DROID datasets and 50 demonstrations collected in the same environment as $\mathcal{D}_{\text{target}}$ (Kitchen-DROID). During the evaluation, we randomize the object pose within a $20\times 20cm$ grid with varying orientations. We report partial (pick) and full success rates (pick and place) by assigning scores of $0.5$ and $1.0$, respectively, and normalize the summed scores by the number of trials resulting in a range of $0-100$.
@@ 5.2 Experimental Evaluation
STRAP (CLIP, $K{=}100$) $\underline{86.0\%\pm 4.1}$ $90.7\%\pm 2.2$ $\underline{42.0\%\pm 0.9}$ $54.7\%\pm 3.3$ $83.3\%\pm 3.0$ $44.9\%\pm 32.7$
STRAP (DINOv2, $K{=}100$) $85.3\%\pm 2.2$ $\underline{91.3\%\pm 2.2}$ $\bm{42.7\%\pm 7.2}$ $57.3\%\pm 7.7$ $85.3\%\pm 2.8$ $45.6\%\pm 32.6$
@@
Does sub-trajectory retrieval improve performance in few-shot imitation learning? STRAP outperforms the retrieval baselines BR and FR on average by $+24.7\%$ and $+25.0\%$ across all 10 tasks (Tab. 1). These results demonstrate the policy’s robustness to varying object poses. BC represents a strong baseline on the LIBERO task as the benchmark’s difficulty comes from pose variations during evaluation. By memorizing the demonstrations, BC achieves high success rates, outperforming BR and FR by $+4.5\%$ and $+4.8\%$ across all 10 tasks. In our real-world experiments, BC performs much worse due to the increased randomization during evaluation. The policy replays the demonstrations in $\mathcal{D}_{\text{target}}$, failing to adapt to new object poses.
@@
Pre-training a policy on $\mathcal{D}_{\text{prior}}$ and fine-tuning it on $\mathcal{D}_{\text{target}}$ (FT) emerges as the most competitive baseline underperforming STRAP by only $-6.4\%$. Training a multi-task policy (MT) matches BC on average but shows improvements when $\mathcal{D}_{\text{prior}}$ contains demonstrations or environments overlapping with $\mathcal{D}_{\text{target}}$. Introducing larger randomization to $\mathcal{D}_{\text{prior}}$ and the evaluation as part of the real-world experiments hurts the performance of both methods. Specifically, we find checkpoint selection for FT difficult as it’s easy for the policy to under- or overfit the demonstrations, leading to degenerate behavior or exact replay. We found MT training challenging as the policy Haldar et al. (2024); Nasiriany et al. (2024) sometimes ignores the language instruction and solves a task seen in $\mathcal{D}_{\text{prior}}$ instead of the conditioned $\mathcal{D}_{\text{target}}$. We hypothesize that this behavior emerges due to the data imbalance of $\mathcal{D}_{\text{prior}}$ and $\mathcal{D}_{\text{target}}$. Finally, augmenting $\mathcal{D}_{\text{prior}}$ with 5000 trajectories from the DROID dataset amplifies these challenges leading to an even larger performance gap. In our real-world evaluations, we find STRAP to experience surprising generalization behavior to poses unseen in $\mathcal{D}_{\text{target}}$. The policy further shows recovery behavior, completing the task even when the initial grasp fails and alters the object’s pose. Since STRAP’s policy training stage is independent of the size of $\mathcal{D}_{\text{prior}}$ but the dataset size is determined by hyperparameter $K$, it can naturally deal with adding in larger datasets like DROID maintaining performance.
@@
How important are sub-trajectories for retrieval and how many should be retrieved? To investigate the efficacy of sub-trajectories, we compare sub-trajectory retrieval with S-DTW (STRAP) to retrieving full trajectories with S-DTW (D-T) in Tab. 1. We find sub-trajectory retrieval to improve performance by $+4.1\%$ across all 10 tasks. We hypothesize that full trajectories can contain segments irrelevant to the task, effectively hurting performance and reducing the accuracy of the cumulative cost. Varying the number of retrieved segments $K$, we find that the optimal value for $K$ is highly task-dependent with some tasks benefiting from retrieving less and some from retrieving more data. We hypothesize that $K$ depends on whether tasks leverage (positive transfer) or suffer (negative transfer) from multi-task training. We report the results for the best $K$ in Tab. 1 and provide the full search in Tab. 9.
@@
How effective are the representations from vision-foundation models for retrieval? Next, we ablate the choice of foundation model representation in STRAP with fixed $K=100$. We compare CLIP, a model trained through supervised learning on image-text pairs, with DINOv2, a self-supervised model trained on unlabeled images. We don’t find any representation to significantly outperform the other with DINOv2 separated from CLIP by only $+0.7\%$ across all 10 tasks. To show the efficacy of vision-foundation models for retrieval, we replace the in-domain feature extractors from prior work (BR, FR) trained on $\mathcal{D}_{\text{prior}}$ with an off-the-shelf DINOv2 encoder model (D-S). Comparing them in their natural configuration, i.e., state-based retrieval using cosine similarity allows for a side-by-side comparison of the representations. Tab. 1 shows the choice of representation to depend on the task with no method outperforming the others on all tasks. Since D-S has no notion of dynamics and task semantics due to single-state retrieval, BR and FR outperform it by $+5.0\%$ and $+4.7\%$, respectively. We highlight that vision foundation models don’t have to be trained on $\mathcal{D}_{\text{prior}}$ and scale much better with increasing amounts of trajectory data and on unseen tasks.

### v2

@@ 3.1 Dynamic Time Warping
A cumulative distance matrix $D$ is computed via dynamic programming as follows: $D(n,1)=\sum_{k=1}^{n}C(k,1)\text{ for }n\in[1:N]\text{ and }D(1,m)=\sum_{k=1}^{m}C(1,k)\text{ for }m\in[1:M]$. Then the following dynamic programming calculation is performed:
@@
where $C(x_{i},y_{j})$ is the distance between points $x_{i}$ and $y_{j}$. We assume this cost matrix is pre-provided, and we describe how we compute this from raw camera images in Sec. 4.3. The optimal alignment between the sequences is found by backtracking from $D(n,m)$ to $D(1,1)$. This guarantees that the start is matched to the start and the end is matched to the end or that the pairs $(x_{0},y_{0})$ and $(x_{n},y_{m})$ are the start and end of the path. This optimal paring path consists of the best possible alignment between $X$ and $Y$ such that the cumulative cost between all matched pairs is minimized. DTW, as described, is widely used in time-series analysis, speech recognition, and other domains where temporal variations exist between sequences. In the context of our retrieval problem, DTW is used to go beyond retrieving exactly matched sequences to matching variable length subsequences, as we describe below.
@@ 4.1 Problem Setting: Retrieval-augmented Policy Learning
Given $\mathcal{D}_{\text{prior}}$ and $\mathcal{D}_{\text{target}}$, the goal is to learn a language-conditioned policy $\pi_{\theta}(a|s,l)$ that can predict optimal actions $a$ in the target environment when prompted with the current state $s$ and language instruction $l$. Assuming we can obtain a measure of success (such as task completion), and a broad set of initial conditions $s_{0}\sim\rho_{\text{test}}(s_{0})$ in the test environment. The objective of policy learning is to determine the policy parameters $\theta$ to maximize the expected success metric when evaluated on test conditions, under the policy $\pi_{\theta}$ and test-time environment dynamics. Since we are only provided a limited corpus of data, $\mathcal{D}_{\text{target}}$, in the target domain, these policy parameters cannot be learned by simply performing maximum likelihood on $\mathcal{D}_{\text{target}}$. Instead, we will present an approach where a smaller, “relevant” subset of the offline dataset $\mathcal{D}_{\text{retrieval}}$ $\subseteq$ $\mathcal{D}_{\text{prior}}$ is retrieved non-parametrically and then mixed with the smaller in-domain dataset $\mathcal{D}_{\text{target}}$ to construct a larger, augmented training dataset, i.e., $\mathcal{D}_{\text{target}}\cup\mathcal{D}_{\text{retrieval}}$, which is most relevant to the desired test-time conditions $\rho_{\text{test}}(s_{0})$. This can then be used for training policies via imitation learning, as we will describe in Sec. 4.5. Doing so avoids an expensive generalist training procedure and rather focuses the learned model to being a high-performing specialist in a particular setting of interest. The key questions becomes - How can we define what subset of the offline dataset $\mathcal{D}_{\text{prior}}$ is relevant to construct $\mathcal{D}_{\text{retrieval}}$?
@@ 4.3 Foundation Model-Driven Relevance Metrics for Retrieval
In this work, we will adopt the insight that vision(-language) foundation models Oquab et al. (2023); Radford et al. (2021) offer off-the-shelf solutions to this problem of measuring the semantic and visual similarities between sub-trajectories, capturing object- and task-centric affordances, while being robust to low-level variations in scene appearance. Trained on web-scale real-world image(-text) data, these models are typically robust to low-level perceptual variations, while providing semantically rich representations that naturally capture a notion of object-ness and semantic correspondence. Denoting a vision foundation model as $\mathcal{F}(\cdot)$, we can compute the pairwise distance of two camera views with an L2 norm [^2] Other cost metrics such as (1-cosine similarity) could be used here as well. in embedding space, i.e., $\mathcal{F}(o_{i})-\mathcal{F}(o_{j})||_{2}$. While aggregation methods such as temporal averaging could be used to go from embedding of a single image to that of a sub-trajectory, they lose out on the actions and dynamics. We instead opt for a sub-trajectory matching procedure based on the idea of DTW Giorgino (2009) and use the embeddings for finding maximally relevant sub-trajectories. Given two sub-trajectories, $\bm{t}_{i}$ and $\bm{t}_{j}$, we compute a pairwise cost matrix $C\in\mathbb{R}^{|\bm{t}_{i}|\times|\bm{t}_{j}|}$, where its value is as computed by:
[…]
@@
[…]
@@
8:  $\mathcal{M}\leftarrow[];$ $\triangleright$ Initialize empty match storage
@@
10:   $D\leftarrow\texttt{computeCostMatrix}(\mathcal{E}_{\text{target}},\mathcal{E}_{\text{prior}});$ $\triangleright$ Eq. 2
11:   $\mathcal{M}_{i,j}\leftarrow\texttt{extractSubTrajectory}(D,T_{\text{prior}});$ $\triangleright$ Dynamic Programming
@@
[…]
@@
17:  sample $\mathcal{B}\sim\mathcal{D}_{\text{target}}\cup\mathcal{D}_{\text{retrieval}}$ to update policy $\pi_{\theta}$ with loss $\mathcal{L}(\mathcal{B};\theta)$ $\triangleright$ Eq. 3
@@ 4.5 Putting it all together: STRAP
To start the retrieval process, we encode image observations in $\mathcal{D}_{\text{target}}$ and $\mathcal{D}_{\text{prior}}$ using a vision foundation model, e.g., DINOv2 Oquab et al. (2023) or CLIP Radford et al. (2021). To best leverage the multi-task trajectories in $\mathcal{D}_{\text{prior}}$, we split the demonstrations in $\mathcal{D}_{\text{target}}$ into atomic chunks based on a low-level motion heuristic. Then we generate matches between chunked $\mathcal{D}_{\text{target}}$ and $\mathcal{D}_{\text{prior}}$ and construct $\mathcal{D}_{\text{retrieval}}$ by selecting the top $K$ matches uniformly across all chunks. To avoid degenerate matches, we limit the step size of the S-DTW backtracking to $[1,2]$ (c.f. Müller (2021)). Combining $\mathcal{D}_{\text{retrieval}}$ with $\mathcal{D}_{\text{target}}$ forms our dataset for learning a policy. In a standard policy learning setting, noisy retrieval data can lead to negative transfer, e.g., when observations similar to the target data are labeled with actions that achieve a different task. Without conditioning, such contaminated samples hurt the policy’s downstream performance. We propose to use a language-conditioned policy to deal with this inconsistency. With conditioning, the policy can distinguish between samples from different tasks, separating misleading from expert actions while benefiting from positive transfer from the additional training data and context of the language conditioning.
@@ 5.1 Experimental Setup
  • DROID-Kitchen: Scaling STRAP to more realistic scenarios, we evaluate STRAP on vegetable pick-and-place in three real kitchen environments. We collect 150 multi-task demonstrations across scenes in $\mathcal{D}_{\text{prior}}$ (Kitchen), with 50 per scene. Task combinations are sampled randomly from three possible tasks per environment and without replacement. Every demonstration consists of two unique tasks with randomized object poses and appearances. $\mathcal{D}_{\text{target}}$ is specific to each environment and contains 3 demonstrations of one of the three possible tasks. To investigate STRAP’s scalability to larger datasets, we construct an additional $\mathcal{D}_{\text{prior}}$ consisting of 5000 demonstrations from the DROID datasets and 50 demonstrations collected in the same environment as $\mathcal{D}_{\text{target}}$ (Kitchen-DROID). During the evaluation, we randomize the object pose within a $20\times 20cm$ grid with varying orientations. We report partial (pick) and full success rates (pick and place) by assigning scores of $0.5$ and $1.0$, respectively, and normalize the summed scores by the number of trials resulting in a range of 01000-100.
@@ 5.2 Experimental Evaluation
STRAP (CLIP, K=100K{=}100) $\underline{86.0\%\pm 4.1}$ $90.7\%\pm 2.2$ $\underline{42.0\%\pm 0.9}$ $54.7\%\pm 3.3$ $83.3\%\pm 3.0$ $44.9\%\pm 32.7$
STRAP (DINOv2, K=100K{=}100) $85.3\%\pm 2.2$ $\underline{91.3\%\pm 2.2}$ $\bm{42.7\%\pm 7.2}$ $57.3\%\pm 7.7$ $85.3\%\pm 2.8$ $45.6\%\pm 32.6$
@@
Does sub-trajectory retrieval improve performance in few-shot imitation learning? STRAP outperforms the retrieval baselines BR and FR on average by $24.7\%$ and $25.0\%$ across all 10 tasks (Tab. 1). These results demonstrate the policy’s robustness to varying object poses. BC represents a strong baseline on the LIBERO task as the benchmark’s difficulty comes from pose variations during evaluation. By memorizing the demonstrations, BC achieves high success rates, outperforming BR and FR by $4.5\%$ and $4.8\%$ across all 10 tasks. In our real-world experiments, BC performs much worse due to the increased randomization during evaluation. The policy replays the demonstrations in $\mathcal{D}_{\text{target}}$, failing to adapt to new object poses.
@@
Pre-training a policy on $\mathcal{D}_{\text{prior}}$ and fine-tuning it on $\mathcal{D}_{\text{target}}$ (FT) emerges as the most competitive baseline underperforming STRAP by only $6.4\%$. Training a multi-task policy (MT) matches BC on average but shows improvements when $\mathcal{D}_{\text{prior}}$ contains demonstrations or environments overlapping with $\mathcal{D}_{\text{target}}$. Introducing larger randomization to $\mathcal{D}_{\text{prior}}$ and the evaluation as part of the real-world experiments hurts the performance of both methods. Specifically, we find checkpoint selection for FT difficult as it’s easy for the policy to under- or overfit the demonstrations, leading to degenerate behavior or exact replay. We found MT training challenging as the policy Haldar et al. (2024); Nasiriany et al. (2024) sometimes ignores the language instruction and solves a task seen in $\mathcal{D}_{\text{prior}}$ instead of the conditioned $\mathcal{D}_{\text{target}}$. We hypothesize that this behavior emerges due to the data imbalance of $\mathcal{D}_{\text{prior}}$ and $\mathcal{D}_{\text{target}}$. Finally, augmenting $\mathcal{D}_{\text{prior}}$ with 5000 trajectories from the DROID dataset amplifies these challenges leading to an even larger performance gap. In our real-world evaluations, we find STRAP to experience surprising generalization behavior to poses unseen in $\mathcal{D}_{\text{target}}$. The policy further shows recovery behavior, completing the task even when the initial grasp fails and alters the object’s pose. Since STRAP’s policy training stage is independent of the size of $\mathcal{D}_{\text{prior}}$ but the dataset size is determined by hyperparameter $K$, it can naturally deal with adding in larger datasets like DROID maintaining performance.
@@
How important are sub-trajectories for retrieval and how many should be retrieved? To investigate the efficacy of sub-trajectories, we compare sub-trajectory retrieval with S-DTW (STRAP) to retrieving full trajectories with S-DTW (D-T) in Tab. 1. We find sub-trajectory retrieval to improve performance by $4.1\%$ across all 10 tasks. We hypothesize that full trajectories can contain segments irrelevant to the task, effectively hurting performance and reducing the accuracy of the cumulative cost. Varying the number of retrieved segments $K$, we find that the optimal value for $K$ is highly task-dependent with some tasks benefiting from retrieving less and some from retrieving more data. We hypothesize that $K$ depends on whether tasks leverage (positive transfer) or suffer (negative transfer) from multi-task training. We report the results for the best $K$ in Tab. 1 and provide the full search in Tab. 9.
@@
How effective are the representations from vision-foundation models for retrieval? Next, we ablate the choice of foundation model representation in STRAP with fixed $K=100$. We compare CLIP, a model trained through supervised learning on image-text pairs, with DINOv2, a self-supervised model trained on unlabeled images. We don’t find any representation to significantly outperform the other with DINOv2 separated from CLIP by only $0.7\%$ across all 10 tasks. To show the efficacy of vision-foundation models for retrieval, we replace the in-domain feature extractors from prior work (BR, FR) trained on $\mathcal{D}_{\text{prior}}$ with an off-the-shelf DINOv2 encoder model (D-S). Comparing them in their natural configuration, i.e., state-based retrieval using cosine similarity allows for a side-by-side comparison of the representations. Tab. 1 shows the choice of representation to depend on the task with no method outperforming the others on all tasks. Since D-S has no notion of dynamics and task semantics due to single-state retrieval, BR and FR outperform it by $5.0\%$ and $4.7\%$, respectively. We highlight that vision foundation models don’t have to be trained on $\mathcal{D}_{\text{prior}}$ and scale much better with increasing amounts of trajectory data and on unseen tasks.